OpenStudy (anonymous):

Algebra 2 Honors Project help please! I need to get this done! D:

OpenStudy (anonymous):

1.Determine how many, what type, and find the roots for f(x) = x^4 + 21x^2 – 100. 2.Determine how many, what type, and find the roots for f(x) = x^3 - 5x^2 – 25x + 125. 3.The following graph shows a seventh-degree polynomial: graph of a polynomial that touches the x axis at negative 5, crosses the x axis at negative 1, crosses the y axis at negative 2, crosses the x axis at 4, and crosses the x axis at 7. Part 1: List the polynomial’s zeroes with possible multiplicities. Part 2: Write a possible factored form of the seventh degree function. 4.Without plotting any points other than intercepts, draw a possible graph of the following polynomial: f(x) = (x + 8)^3(x + 6)^2(x + 2)(x – 1)^3(x – 3)^4(x – 6).

OpenStudy (anonymous):

@Kawaiicat123 Can you help me please?? ;A;

OpenStudy (anonymous):

Four solutions were found : x = 5 x = -5 x = 0.0000 - 2.0000 i x = 0.0000 + 2.0000 i Reformatting the input : Tiger made changes, which won't affect the solution, to your input: (1): "x2" was replaced by "x^2". 1 more similar replacement(s). Rearrange : Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : 0-(x^4-21*x^2-100)=0 Step by step solution : Step 1 : Simplify 0 - x4-21x2-100 Trying to factor by splitting the middle term 1.1 Factoring -x4+21x2+100 The first term is, -x4 its coefficient is -1 The middle term is, +21x2 its coefficient is 21 The last term, "the constant", is +100 Step-1 : Multiply the coefficient of the first term by the constant -1 • 100 = -100 Step-2 : Find two factors of -100 whose sum equals the coefficient of the middle term, which is 21 -100 + 1 = -99 -50 + 2 = -48 -25 + 4 = -21 -20 + 5 = -15 -10 + 10 = 0 -5 + 20 = 15 -4 + 25 = 21 Tiger found it!! Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, (-4) and (25) -x4 - 4x2 + 25x2 + 100 Step-4 : Add up the first 2 terms, pulling out like factors : -x2 • (x2+4) Add up the last 2 terms, pulling out common factors : 25 • (x2+4) Now add up the four terms of step 3 : (-x2+25) • (x2+4) Which is the desired factorization

OpenStudy (anonymous):

Four solutions were found : x = 5 x = -5 x = 0.0000 - 2.0000 i x = 0.0000 + 2.0000 i Reformatting the input : Tiger made changes, which won't affect the solution, to your input: (1): "x2" was replaced by "x^2". 1 more similar replacement(s). Rearrange : Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : 0-(x^4-21*x^2-100)=0 Step by step solution : Step 1 : Simplify 0 - x4-21x2-100 Trying to factor by splitting the middle term 1.1 Factoring -x4+21x2+100 The first term is, -x4 its coefficient is -1 The middle term is, +21x2 its coefficient is 21 The last term, "the constant", is +100 Step-1 : Multiply the coefficient of the first term by the constant -1 • 100 = -100 Step-2 : Find two factors of -100 whose sum equals the coefficient of the middle term, which is 21 -100 + 1 = -99 -50 + 2 = -48 -25 + 4 = -21 -20 + 5 = -15 -10 + 10 = 0 -5 + 20 = 15 -4 + 25 = 21 Tiger found it!! Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, (-4) and (25) -x4 - 4x2 + 25x2 + 100 Step-4 : Add up the first 2 terms, pulling out like factors : -x2 • (x2+4) Add up the last 2 terms, pulling out common factors : 25 • (x2+4) Now add up the four terms of step 3 : (-x2+25) • (x2+4) Which is the desired factorization

OpenStudy (anonymous):

Whoah, holy poop, what is that to???